# A Multi-Perspective Analysis of Memorization in Large Language Models

### v3

@@ Abstract
The memorization behavior of Large Language Models (LLMs) means that LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this study, we comprehensively discussed memorization from various perspectives and extended the discussion to less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed the dynamics of how unmemorized sentences transition to memorized sentences. (2) Through input analysis, we found the frequency of tokens related to memorization and discovered the boundary effect when generating memorized or unmemorized sequences. (3) Through an analysis of embedding and decoding, we showed the LLM can generate paraphrased memorized sentences and sentences of different memorization scores clustered in the embedding space. The entropy also shows a boundary effect when generating memorized or unmemorized content. (4)We trained a Transformer model to predict the memorization of different models, showing the possibility of predicting memorization and paraphrase memorization. Further analysis shows that the prediction of unmemorized sequences is easier than the prediction of memorized content, which can be explained by the significance of the boundary effect.
@@ 1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the recent strongest GPT-4 OpenAI et al. (2024), have not just revolutionized the Natural Language Processing (NLP) but nearly the whole artificial intelligence research field. Though their performances on various tasks are surprising, we still know little about the underlying mechanism of Large Language Models whose structures are based on neural models known as black-box Alain and Bengio (2018). Due to the enormous model size and pre-train data size, unprecedented phenomena in previous machine learning are also identified in the LLM, while in this research, we discuss the memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM sometimes will verbatim output the same sentence recorded in their pre-training corpus. Memorization is a coin with two sides. On the one hand, utilizing such memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple uses. On the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Though such a phenomenon has been utilized or studied since the merge of LLMs, detailed analysis of memorization is still under-explored.
In this research, we discuss memorization in detail and analyze it from several perspectives, e.g., the pre-train data level, the encoding level, the generation level, and finally, we try to predict memorization. Following those studies, we found that:
@@
  • For both memorization and unmemorized sentences, even though their numbers increase/decrease with the model size, the trend is not linear, showing the maximum capacity to memorize sentences exists. The number of memorized sentences decreases sub-linearly with the continuation size while increasing above linearly with the context size. We also study how unmemorized sentences transition to memorized sentences and vice versa. We found that memorized sentences are more likely to be memorized when scaling the model size, while low-memorized sentences will be gradually memorized.
  • Through a study of the encoding dynamics, we found that sentences with different memorization scores are clustered in the embedding space. Additionally, with the generation of tokens, the embedding similarly remains basically stable while the Euclidean distance is shrinking. Moreover, the larger the model, the more significant the difference in the Cosine Similarity. This may help to explain why the larger model performs better than the smaller-size model.
  • In both the n-gram and decoding entropy dynamics, we found that a boundary phenomenon exists when the model generates memorized or unmemorized sentences. In the n-gram, the frequency when generating unmemorized tokens decreases while increasing for memorized tokens. In the entropy analysis, the entropy suddenly increases when unmemorized sentences are generated, while it decreases for memorized sentences.
  • Finally, we trained a naive Transformer model to predict memorization based on context. We show that prediction accuracy over tokens is high and grows with the model size, while it is still hard to predict correctly for a whole sentence
@@ 2.1 Scaling Laws of LLM
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how LLM Scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how LLM scaling affects both translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size such as digital addition Nye et al. (2021), reasoning Lampinen et al. (2022). Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
In the field of memorization, Carlini et al. (2023) discussed the memorization across model size and found that the number of memorized texts grows with the model size and context size. Additionally, Biderman et al. (2023a) also discussed similar topics and found that a large portion of memorized text in a small-size model is also memorized by a larger model, showing that memorized texts may share certain common features.
@@ 2.2 Memorization
Before LLM, a phenomenon close to memorization in real-world practices is known as over-fitting Tetko et al. (1995), which means a near-zero loss in the train set while bad performance in the test set. Zhang et al. (2017) described that the neural model could reach near-zero loss on the training set, which suggests the neural model perfectly memorized the relation between input and data label. However, memorization is different from over-fitting as LLM has good performance in the test set or the real world. Feldman (2020) has conducted a theoretical analysis explaining why memorization happens and is actually needed. They showed in a long-tail distribution where many categories only have very few samples. As the neural model cannot extract general features from the limited number of samples in those categories, the best choice for the neural model is just to memorize them and compare them with data in the test set.
In the era of LLM, different from classification models, the LLM can directly output their pre-training content that can be used from a knowledge graph Petroni et al. (2019); AlKhamissi et al. (2022) while also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Regarding specific research on memorization behavior, Tirumala et al. (2022) discussed training dynamics of memorization, showing large models tend to memorize training samples more easily. Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, suggesting they are inter-correlated. Biderman et al. (2023a) also discussed memorization in LLM regarding their training phase and the overlap between different model sizes, showing the possibility of using smaller LLM to predict memorization of larger LLM.
Compared to the previous memorization studies, our research discussed memorized text and extended this scope to less and unmemorized text. Additionally, we also discussed both the encoding and decoding dynamics of generations. Finally, we trained a Transformer model to predict memorization and analyze the results.
Figure 1: Memorization and Research Scope
@@ 3.1.1 Memorization Criteria
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the true continuation tokens. $L(Y)$ means the length of the true continuation tokens. If $M(X,Y)==1$, this sequence is completely memorized under the context sequence, which is called K-extractable. Similarly, we also define the sequence $Y$ as completely not memorized if $M(X,y)==0$. Since the memorization score has high granularity, in some experiments, we also classify sentences into several memorization levels based on their memorization scores.
@@ 3.1.2 Prediction Criteria
In this study, we use two criteria to evaluate model performance in predicting memorization, e.g., Token-Level Accuracy and Full Accuracy.
We define the prediction for a sequence of generated tokens as $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where is $\hat{y}$ is the golden binary label.
@@
The Full Accuracy means that the prediction of token level at a sentence must be entirely correct, which means $T(\hat{X},\hat{Y})$ must equal $1$.
@@ 4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into ten sets with a memorization difference of 0.1 The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
  • We can see that the number of sentences with low memorization scores (0-0.3) is significantly higher than sentences with high memorization scores. This shows that though memorization exists in the LLM, most of the pre-train data are still not memorized.
  • Among sentences with high memorization scores, the fully memorized sentence increases faster than others. This indicates that LLM tends to fully memorize sentences rather than memorizing a piece of it. We can also see that the number of sentences with low memorization scores (0, 0.1, etc) decreases with the increase in model size. This shows that unmemorized sentences will gradually memorized with the increase in model size.
  • Additionally, we can see that the sentences that are being memorized or memorized do not increase or decrease linearly with the increase in model size. For fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from this figure for the unmemorized sentence. This shows an upper limit exists for memorization that the LLM cannot memorize the whole corpora even with enough model size.
4.1.2 The Factor of Context Size and Complement Size
We change the number of prompted tokens while fixing the complement size and then observe the memorization. From the Figure 2 (b) and (c), we can tell that:
  • The decrease in memorized sentences with the complement size is not linear. When we increase the complement size from 64 to 96, the number of memorized sentences does not change too much compared to the change when the complement size changes from 32 to 48. This also suggests some sentences are firmly memorized.
  • Additionally, the drop in the number of memorized sentences with the increased complement size is more significant in the larger model. This shows that though larger models memorize more sentences, the memorized sentence is not as firm as sentences memorized by small models.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to a near-exponential increase in the number of memorized sentences. Additionally, the increase is more significant in the larger model, showing that more potential memorized contents can be elicited by giving more context compared to the small model.
@@ 4.1.2 The Factor of Context Size and Complement Size
4.2 Memorization Transition Across Model Size
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We classify sentences with memorization scores with a range difference of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model.
From the Figure 3, we can conclude:
  • Most sentences will stay at their previous state even when trained in a larger model. We can tell this by looking at the diagonal of those two transition matrices. Additionally, the higher the memorization score, the harder the sentence will transition to another state. For highly memorized sentences, more than $90\%$ of sentences will remain being memorized compared to low memorized sentences.
  • We can also see with the increase in the model size, the sentence is more inclined to stay in its original state. In the transition from the 2.8b model to the 12b model, we can tell that the transition probability at the diagonal is higher than the 410m to 2.8b transition matrix counterpart. This suggests that the memorized or unmemorized sentences tend to be fixed with the increase in model size.
  • We also see that even for a highly memorized sentence, there is still a little chance that the sentence will transit to a low memorized state, suggesting some sentences may be memorized randomly rather than following a specific feature. Such randomness makes it possible for them to transit to a low memorized state when trained in a large model.
@@ 4.3.1 Token Level Frequency Analysis
  • We can observe a very obvious boundary effect. At the index around 32, which is the first generated token, the frequency drops and rises for the memorized sentence (positive boundary effect) while rises and drops for the unmemorized sentences (negative boundary effect). Additionally, the negative boundary effect is more obvious than the positive boundary effect.
  • For the half-memorized sentence, we see the frequency increase and drop around the index of 39, which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized, and the later half is mostly unmemorized, meaning that the memorized tokens are distributed in a near continuous way rather than scattered in the generated tokens. We also notice the repetitive pattern of half-memorized sentences and unmemorized sentences, which is the negative boundary effect around the index of 39.
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having a higher frequency, showing the frequency of the initial tokens is more important than the following tokens, e.g., remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results: the low frequency of the first several tokens makes the following sentence easy to forget.
@@ 4.3.2 Sentence Level Frequency Analysis
We have shown the existence of a boundary effect in the above experiment. In this section, we discuss how such boundaries exist across model sizes and whether there are other factors influencing memorization. We calculate the average frequency of context, continuation, and also the frequency difference at the boundary in Table 1. From this table, we can obtain the following:
  • First, we can see that the frequency of one-gram is significantly higher than the frequency of two-gram. In the context and continuation, the average frequency is roughly 3.5 times higher than the two-gram frequency. Additionally, such boundary effect is consistent in the two-gram setting. However, the positive boundary effect is less obvious due to the decrease of frequency when computed by 2-gram. However, as the unit is billion, the actual frequency gap is still a million level.
  • In the context and continuation, in memorized sentences, the frequency is higher in lower in the context but higher in the continuation, while the unmemorized sentence shows the opposite. This suggests the boundary effect also exists in a broader sequence level, though less obvious.
  • However, the half-memorized sentence shows different results in that the frequency of continuation tokens is higher than the average frequency of context, which is caused by the increase of frequency before reaching the first generated unmemorized tokens indicated in Figure 4.
  • In the specific boundary frequency difference around the first generated token, we can see a decreasing trend for both positive and negative boundary effects. However, the decrease in the positive boundary effect makes it less significant, while the decrease in the negative boundary effect makes it more significant. These results suggest that the significance of the positive boundary corresponds to how easily the sentence can be memorized. In contrast, the significance of the negative boundary effect corresponds to how easily the sentence can not be memorized.

### v4

@@ Abstract
Memorization, meaning that Large Language Models (LLMs) can generate content used to train them, is one of the special behaviors of LLMs. Through studies by previous research, little attention has been paid to what makes sentences memorized, how model size affects it, and the dynamics of generating it. In this study, we discuss memorization from multiple perspectives, including scaling model size, input and output dynamics, and less unmemorized content, and reveal: (1) The inter-correlation between memorized/unmemorized sentences, model size, continuation size, and context size, as well as the transition dynamics between sentences of different memorization scores. (2) A boundary effect when generating memorized/unmemorized content and its relation to model size. (3) Sentences of different memorization scores cluster in the embedding space are observed through an embedding dynamics analysis. Entropy analysis in decoding also shows a boundary effect when generating memorized/unmemorized content, albeit oppositely. (4) The possibility of predicting memorization and its relation to model size and continuation length. Further analysis showed that unmemorized sequences are easier to predict than memorized content, which can be explained by the significance of the boundary effect.
@@ 1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the GPT-4 OpenAI et al. (2024), have revolutionized the Natural Language Processing (NLP) and the artificial intelligence research field. Though surprised by their performance on various tasks, we still know little about the underlying mechanism of LLMs whose structures are based on neural models known as black-box Alain and Bengio (2018). Additionally, due to its unprecedented model size and pre-train data size, LLMs start to present unique behaviors Wei et al. (2022), and one of them is memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM can generate the same content recorded in their pre-train corpus under certain contexts. On the one hand, utilizing memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple uses. On the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Previous research Tirumala et al. (2022); Carlini et al. (2023); Biderman et al. (2023b) have studied memorization at the macro level, leaving more micro yet important questions left under-explored, e.g., what makes sentences memorized, what role does the model size play, the input and output dynamics while generating memorized or unmemorized content and how it is related to less memorized content. To answer the above questions, we conducted a multi-perspective study relating memorization to factors like the scaling model size, the input and output dynamics, general statistics, and the prediction of memorization, and we found that:
  1. (I)
@@
    For both memorized and unmemorized sentences, the increase or decrease in model size follows a non-linear trend, indicating a maximum capacity for memorization. The number of memorized sentences decreases sub-linearly with continuation size and increases super-linearly with context size. Moreover, transitions between sentences of different memorization scores were observed, showing that only limited sentences will be memorized with the increase in model size while most sentences remain unmemorized.
  2. (II)
    We conducted an input dynamics study focusing on the n-gram frequency of inputs. A boundary effect was observed when the model began generating memorized/unmemorized content: the frequency of generating unmemorized tokens suddenly decreased while increased for memorized tokens. This boundary effect is also observed at the sentence level and is more obvious in smaller models, indicating that the boundary effect’s significance determines the ease of memorization for a sentence.
  3. (III)
    Additionally, we analyzed decoding dynamics by examining entropy over vocabulary and the drift of decoded embeddings. Entropy analysis revealed an inverse boundary effect, where entropy suddenly increases for unmemorized content and decreases for memorized content. The embedding dynamics analysis showed sentences with different memorization scores cluster in the embedding space, where the mutual embedding distance grows with model size The close distance between sentences with close memorization scores also suggests the existence of paraphrased memorization.
  4. (IV)
    Finally, we trained a Transformer model to predict memorization based on context, showing the possibility of predicting memorization. Additionally, the results suggest that predicting memorization is easier in large models and easier when predicting unmemorized content, which can be explained by the significance of the boundary effect.
@@ 2.1 Scaling Laws of LLM
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how the scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how scaling affects translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size. Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
Regarding scaling in the field of memorization, Carlini et al. (2023) discussed the memorization across model size and found that the number of memorized texts grows with the model size and context size. Additionally, Biderman et al. (2023a) also discussed similar topics and found that a large portion of memorized text in a small-size model is also memorized by a larger model, showing that memorized texts may share certain common features.
@@ 2.2 Memorization
Before LLM and in real-world practices, over-fitting is close to memorization Tetko et al. (1995), which means a near-zero loss in the train set that suggests the neural model perfectly memorized the relation between input and data label Zhang et al. (2017). However, memorization is different from over-fitting as LLM has good performance in the real world while overfitting is usually accompanied by low performance in the test set. Feldman (2020) has conducted a theoretical analysis in classification models explaining why memorization is actually needed. They showed that in a long-tail distribution where many categories only have a few samples, the neural model cannot extract general features from them. The best choice for the neural model is just to memorize them and compare them with data in the test set.
Figure 1: Memorization and Research Scope in this study
For LLMs, different from classification models, they directly generate their pre-train content making the memorization directly observable. which in one way can be used to form a knowledge graph Petroni et al. (2019); AlKhamissi et al. (2022) but in another way also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Previous research has discussed memorization from a macro level, Tirumala et al. (2022) showing large models tend to memorize samples more easily in the training phase. Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, showing their inter-correlations. Biderman et al. (2023a) studied memorization in LLM regarding their training phase and the overlap between different model sizes, showing the existence of commonly memorized sentences.
Compared to the previous memorization studies, our research focuses on more micro and under-explored questions, e.g., the dynamics while generating them, the reason why some sentences are memorized, and the prediction of memorization.
@@ 3.1.1 Memorization Criteria
[…]
@@ 3.1.2 Prediction Criteria
This study uses two criteria to evaluate model performance in predicting memorization, e.g., Token-Level Accuracy and Full Accuracy.
The prediction for a sequence of generated tokens is $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where each $\hat{y}$ is the golden label.
@@
The Full Accuracy means that the prediction of token level at a sentence must be entirely correct, which means $T(\hat{X},\hat{Y})==1$.
@@ 4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into ten sets with a memorization difference of 0.1. The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
  1. (I)
    The number of sentences with low memorization scores (0-0.3) is significantly higher than those with high memorization scores, indicating that most of the pre-train data are not memorized despite the existence of memorization in LLMs.
  2. (II)
    Among sentences with high memorization scores, the count of fully memorized sentences increases more rapidly, suggesting that LLMs have a propensity to fully memorize sentences rather than partially. Additionally, the number of sentences with low memorization scores (0, 0.1) decreases as the model size increases, indicating that unmemorized sentences gradually become memorized with larger models.
  3. (III)
    The increase or decrease in the number of memorized or unmemorized sentences is not linear with respect to model size. There is a noticeable increase in numbers for fully memorized sentences from 70 million to 2.8 billion parameters, compared to 2.8 billion to 12 billion parameters. A similar decreasing trend for unmemorized sentences is observed. This suggests a capacity for memorization, implying LLMs cannot memorize the entire corpus even with sufficiently large model sizes.
4.1.2 Context and Complement Size
We change the number of prompted tokens while fixing the complement or context size and then observe the number of fully memorized sequences. From the Figure 2 (b) and (c), we can tell that:
  1. (I)
    The decrease in the number of memorized sentences with increasing complement size is not linear. For instance, the complement size increase from 64 to 96 results in a relatively minor decrease compared to the change from 32 to 48, indicating some sentences are firmly memorized.
  2. (II)
    The reduction in memorized sentences with increased complement size is more obvious in larger models. This demonstrates that although larger models memorize more sentences, their memorization is less robust compared to smaller models.
  3. (III)
    The increase in memorized sentences with context size is also non-linear, with longer context leading to an almost exponential rise in the number of memorized sentences. This increase is more significant in larger models, indicating that more contents are potentially memorized in large models, which can be elicited by giving longer context.
@@ 4.1.2 Context and Complement Size
4.2 Memorization Transition
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We classify sentences with memorization scores with a range difference of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model, and we can conclude:
  1. (I)
    Most sentences remain in their previous state even when trained with a larger model, as indicated by the diagonal entries in the transition matrices. Additionally, the higher the memorization score, the less likely the sentence will transition to another state. For highly memorized sentences, over 90% remain memorized compared to those with low memorization scores.
  2. (II)
    With increasing model size, sentences are more likely to stay in their original state. For example, the transition probability at the diagonal is higher when moving from the 2.8 billion to the 12 billion parameter model compared to the transition from the 410 million to the 2.8 billion parameter model. This suggests memorized or unmemorized states become more fixed as the model size increases.
  3. (III)
    Even for highly memorized sentences, there is a small chance of transitioning to a low memorized state, implying that some sentences may be memorized randomly rather than due to specific features. This randomness allows for the possibility of transitioning to a low memorized state when trained on a larger model.
@@ 4.3.1 Token Level Frequency Analysis
  1. (I)
    A clear boundary effect is observed. Around index 32, representing the first generated token, the frequency drops and then rises for memorized sentences (positive boundary effect), whereas it rises and then drops for unmemorized sentences (negative boundary effect). The negative boundary effect is more pronounced than the positive one.
  2. (II)
    For the half-memorized sentence, we see the frequency increase and drop ((negative boundary effect)) around the index of 39, which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized, and the later half is mostly unmemorized, meaning that the memorized tokens are distributed in a near continuous way rather than scattered in the generated tokens.
  3. (III)
    The positive boundary effect in memorized sentences suggests that memorization is driven by the higher frequency of initial tokens, implying that remembering the first few tokens makes the entire sentence easier to memorize. Conversely, the negative boundary effect in unmemorized sentences indicates that the low frequency of initial tokens makes the following sequences easier to forget.
@@ 4.3.2 Sentence Level Frequency Analysis
Given the existence of the boundary effect at the token level, we discuss how such an effect exists across model sizes and whether other factors influence memorization at the sequence level. We calculate the average frequency of context and continuation tokens and frequency difference at the boundary in Table 1, and we can obtain:
  1. (I)
    The frequency of one-grams is significantly higher than that of two-grams, approximately 3.5 times higher on average in both context and continuation. The boundary effect is consistent in the two-gram setting, though the positive boundary effect is less obvious due to the frequency drop when computed with two-grams. Despite this, the actual frequency gap remains substantial (million level) when considering the unit is billion.
  2. (II)
    In memorized sentences, the frequency is lower in the context and higher in the continuation, whereas unmemorized sentences exhibit the opposite pattern. This suggests that the boundary effect persists at a broader sequence level, although it is less obvious.
  3. (III)
    For half-memorized sentences, the frequency of continuation tokens is higher than the context average frequency. This is due to the frequency increase before reaching the first generated unmemorized tokens, as indicated in Figure 4.
  4. (IV)
    In the Boundary Frequency Difference column, we see a decreasing trend in both positive and negative boundary effects. However, the decrease in the positive boundary effect makes it less significant, while the decrease in the negative boundary effect makes it more significant. These results imply that the significance of the positive boundary effect correlates with the ease of memorizing a sentence. In contrast, the significance of the negative boundary effect correlates with the ease of not memorizing a sentence.